average age of a boy and a girl is 14 years. The average age of the same boy, the girl and her twin sister is 16 years. What is the boy's age?
Answer by jorel1380(3719)

Let x and y be the boy and the girl, respectively. Then:
x+y/2=14
Also
x+2y/3=16
So:
x+y=28
x+2y=48
y=20
x=8
The boy is 8, and the girl, and her twin sister, is 20. ☺☺☺☺